Tierney Company begins operations on April 1. Information from job cost sheets shows the following.

Manufacturing Costs Assigned

Job
Number


April


May


June

Month
Completed

10 $7,100 $6,400 May
11 5,900 5,900 $3,300 June
12 1,500 April
13 6,500 6,100 June
14 6,800 4,000 Not complete


Job 12 was completed in April. Job 10 was completed in May. Jobs 11 and 13 were completed in June. Each job was sold for 25% above its cost in the month following completion.

(a)

  • Your Answer
  • Correct Answer (Used)

What is the balance in Work in Process Inventory at the end of each month?

Work in Process Inventory
April 30

$

May 31

$

June 30

$

(b)

What is the balance in Finished Goods Inventory at the end of each month?

Finished Goods Inventory
April 30

$

May 31

$

June 30

$

c. what is the gross profit for each month?
Answer #1
a.) Work in Process Inventory Job
Apr-30 $ 13,000 =7100+5900 10 & 11
May-31 $ 25,100 =5900+5900+6500+6800 11,13 & 14
Jun-30 $ 10,800 =6800+4000 14
b.) Finished Goods Inventory Job
Apr-30 $ 1,500 12
May-31 $ 13,500 =7100+6400 10
Jun-30 $ 27,700 =5900+5900+3300+6500+6100 11 & 13
c.) Gross Profit Job
Apr-30                                       -  
May-31 $ 375 =1500*25% 12
Jun-30 $ 3,375 =13500*25% 10
Jul-31 $ 6,925 =27700*25% 11 & 13